# Dusit Niyato, , Bo Ai

### v1

@@ Appendix B Proof of Theorem 1
where $\sum_{\mathbf{m}}{}=\sum_{m_{1}=1}^{M}{\sum_{m_{2}=1}^{M}{\sum_{m_{3}=1}^{M}{\sum_{m_{4}=1}^{M}{}}}}$. Note that different phase-shifts with different $m$-subscript are i.i.d. and many cross-terms equal to 0 due to the existence of phase-shifts. By discussing all possible combinations of $m_{1}\sim m_{4}$, we can compute (43) into closed-form as shown in (10). As for $\mathbb{E}\{\|\mathbf{v}_{k}\|^{2}\}$, we have $\mathbb{E}\{\|\mathbf{v}_{k}\|^{2}\}=\mathbb{E}\{\hat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{W}_{k}\hat{\mathbf{g}}_{k}\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\})=\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\bar{\mathbf{R}}_{k}).$ Finally, combining all results in this appendix, we finish the proof.
@@ Appendix C Proof of Theorem 2
We provide detailed proof steps of Theorem 2 in this appendix. We start from $\mathbb{E}\left\{\mathbf{b}_{kk}\right\}$ and we can easily compute the $m$-th element of it in closed-form as $\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{R}}_{mk})$ based on the results in Appendix B. As for $[\mathbf{\Xi}_{kl}]_{nm}=\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}$, we can compute it in closed-form for all possible AP-UE combinations. For $m\neq n,l\notin\mathcal{P}_{k}$, we have $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\mathbb{E}\{\mathbf{g}_{ml}^{H}\mathbf{v}_{mk}\}\mathbb{E}\{\mathbf{v}_{nk}^{H}\mathbf{g}_{nl}\}=0$. For $m\neq n,l\in\mathcal{P}_{k}$, $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\mathrm{tr}(\mathbf{W}_{mk}\mathbf{B}_{mkl})\mathrm{tr}(\mathbf{W}_{nk}^{H}\mathbf{B}_{nlk})$, where $\mathbf{B}_{mkl}=\mathbb{E}\{\hat{\mathbf{g}}_{mk}\mathbf{g}_{ml}^{H}\}=\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{ml}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$ and $\bar{\mathbf{G}}_{mkk}+\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{ml}$ for $l=k$. For $m=n,l\notin\mathcal{P}_{k}$, $\hat{\mathbf{g}}_{mk}$ and $\mathbf{g}_{ml}$ are independent, so we can derive $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{G}}_{mll}\mathbf{W}_{mk}\bar{\mathbf{G}}_{mkk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\bar{\mathbf{G}}_{mkk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\bar{\mathbf{G}}_{mll}\hat{\mathbf{R}}_{mk}\mathbf{W}_{mk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\hat{\mathbf{R}}_{mk})$, by the method discussed in (41). As for $m=n,l\in\mathcal{P}_{k}$, $\hat{\mathbf{g}}_{mk}$ and $\mathbf{g}_{ml}$ are not independent. Based on the similar method as (42), we can easily compute $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\varphi_{mkl}+\upsilon_{mkl}$. Moreover, we have $\mathbb{E}\{\|\mathbf{v}_{mk}\|^{2}\}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{W}_{mk}\bar{\mathbf{R}}_{mk})$. Based on all derived results, we can easily derive the closed-form results in Theorem 2.
@@ Appendix D Proof of Corollary 1
By applying (32), we can represent $\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}$ in (22) as (44) on the top of this page. By vectorizing $\mathbf{w}_{k}=\mathrm{vec}(\mathbf{W}_{k})\in\mathbb{C}^{M^{2}N^{2}}$, we can derive $\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})=\mathbf{w}_{k}^{H}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})$, applying (34) by letting $\mathbf{A}=\mathbf{W}_{k}$ and $\mathbf{B}=\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\}$, respectively. As for $\mathbb{E}\{\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{W}_{k}\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H})\}$, we can apply (36) to derive $\mathbb{E}\{\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{W}_{k}\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H})\}\!\!=\!\!\mathbf{w}_{k}^{H}\mathbb{E}\{[(\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H})^{T}\otimes(\mathbf{g}_{l}\mathbf{g}_{l}^{H})]\}\mathbf{w}_{k},$ by letting $\mathbf{A}=\mathbf{W}_{k}$, $\mathbf{B}=\mathbf{g}_{l}\mathbf{g}_{l}^{H}$, $\mathbf{C}=\mathbf{W}_{k}$, and $\mathbf{D}^{T}=\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}$, respectively. For $\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\})$, based on (35), we can obtain $\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\})=\mathbf{w}_{k}^{H}(\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\}^{T}\otimes\mathbf{I}_{MN})\mathbf{w}_{k}$, by letting $\mathbf{A}=\mathbf{W}_{k}$, $\mathbf{B}=\mathbf{I}_{MN}$, $\mathbf{C}=\mathbf{W}_{k}$, and $\mathbf{D}^{T}=\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\}$, respectively. Thus, by combining all these terms, we can construct (44) as
$\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}=$ (45)
$\frac{p_{k}|\mathbf{w}_{k}^{H}\mathrm{vec(}\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})|^{2}}{\mathbf{w}_{k}^{H}\left[\begin{array}[]{c}\sum_{l=1}^{K}{p_{l}\mathbb{E}\{[(\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H})^{T}\otimes(\mathbf{g}_{l}\mathbf{g}_{l}^{H})]\}}\\ -p_{k}|\mathbf{w}_{k}^{H}\mathrm{vec(}\mathbb{E}\{\mathbf{g}_{k}\hat{\mathbf{g}}_{k}^{H}\})|^{2}\!+\!(\mathbb{E}\{\hat{\mathbf{g}}_{k}\hat{\mathbf{g}}_{k}^{H}\}^{T}\otimes\mathbf{I}_{MN})\\ \end{array}\right]\!\mathbf{w}_{k}}.$
Note that (45) is a standard Rayleigh quotient with respect to $\mathbf{w}_{k}$. Thus, by applying [5, Lemma B.10], we can derive the optimal $\mathbf{w}_{k}^{*}$, which can maximize (45), as in (24), leading to the maximum SINR value as in (23).
@@ Appendix E Proof of Theorem 3
We first reformulate terms of $\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}}$ given in (8). By applying (34), we have $\mathrm{tr}(\mathbf{W}_{k}^{H}\bar{\mathbf{R}}_{k})=\mathbf{w}_{k}^{H}\bar{\mathbf{r}}_{k}$. Moreover, by applying results in Appendix A, we can easily represent $\mu_{kl}$, $\varepsilon_{kl}$, and $\mathrm{tr}(\mathbf{W}_{k}\bar{\mathbf{R}}_{k}\mathbf{W}_{k}^{H})$ in (8) as $\mu_{kl}=\mathbf{w}_{k}^{H}(\bar{\mathbf{G}}_{kk}^{T}\otimes\bar{\mathbf{G}}_{ll})\mathbf{w}_{k}+\mathbf{w}_{k}^{H}(\bar{\mathbf{G}}_{kk}^{T}\otimes\check{\mathbf{R}}_{l})\mathbf{w}_{k}+\mathbf{w}_{k}^{H}(\hat{\mathbf{R}}_{k}^{T}\otimes\bar{\mathbf{G}}_{ll})\mathbf{w}_{k}+\mathbf{w}_{k}^{H}(\hat{\mathbf{R}}_{k}^{T}\otimes\check{\mathbf{R}}_{l})\mathbf{w}_{k}$ and $\varepsilon_{kl}=p_{k}p_{l}\tau_{p}^{2}|\mathbf{w}_{k}^{H}\tilde{\mathbf{r}}_{lk}|^{2}$, and $\mathrm{tr}(\mathbf{W}_{k}\bar{\mathbf{R}}_{k}\mathbf{W}_{k}^{H})=\mathbf{w}_{k}^{H}(\bar{\mathbf{R}}_{k}^{T}\otimes\mathbf{I}_{MN})\mathbf{w}_{k}$, respectively. As for $\omega_{k}$ in (10), the first three terms can also be easily reformulated as $p_{k}\tau_{p}\mathbf{w}_{k}^{H}\bar{\mathbf{g}}_{kk}\tilde{\mathbf{r}}_{kk}^{H}\mathbf{w}_{k}+p_{k}\tau_{p}\mathbf{w}_{k}^{H}\tilde{\mathbf{r}}_{kk}\bar{\mathbf{g}}_{kk}^{H}\mathbf{w}_{k}-\mathbf{w}_{k}^{H}(\bar{\mathbf{G}}_{kk}^{T}\otimes\bar{\mathbf{G}}_{kk})\mathbf{w}_{k}$. However, as for the last three terms in (10), it is difficult to directly handle them since $\mathbf{W}_{k}$ has been divided into many block sub-matrices. To efficiently handle them, we focus on their original expectation form $\mathbb{E}\{\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\bar{\mathbf{g}}_{k}\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}\bar{\mathbf{g}}_{k}\}$. By applying (36), we have $\mathbb{E}\{\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\bar{\mathbf{g}}_{k}\bar{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}\bar{\mathbf{g}}_{k}\}=\mathbf{w}_{k}^{H}\underset{\mathbf{\Upsilon}_{k}}{\underbrace{\mathbb{E}\{(\bar{\mathbf{g}}_{k}\bar{\mathbf{g}}_{k}^{H})^{T}\otimes\bar{\mathbf{g}}_{k}\bar{\mathbf{g}}_{k}^{H}\}}}\mathbf{w}_{k}$. Then, we focus on the computation of $\mathbf{\Upsilon}_{k}\in\mathbb{C}^{M^{2}N^{2}\times M^{2}N^{2}}$, we define $\tilde{\mathbf{G}}_{k}=\bar{\mathbf{g}}_{k}\bar{\mathbf{g}}_{k}^{H}\in\mathbb{C}^{MN\times MN}$ and the $(d,c)$-th element of $\tilde{\mathbf{G}}_{k}$ is $\tilde{G}_{k,dc}=\bar{g}_{m_{1}k,n_{1}}\bar{g}_{m_{2}k,n_{2}}^{*}e^{j\theta_{m_{1}k}}e^{-j\theta_{m_{2}k}}$, where these variables are defined in Theorem 3. Then, by applying the characteristic of Kronecker-product, we can represent the $(c,d)$-th block sub-matrix of $\mathbf{\Upsilon}_{k}$ as $\mathbf{\Upsilon}_{k,cd}=\mathbb{E}\{\tilde{G}_{k,dc}\tilde{\mathbf{G}}_{k}\}\in\mathbb{C}^{MN\times MN}$. Finally, by plugging $\tilde{G}_{k,dc}$ into $\mathbf{\Upsilon}_{k,cd}$ and applying the features of random phase-shifts, we can easily compute $\mathbf{\Upsilon}_{k}$ in closed-form as $\bar{\mathbf{\Upsilon}}_{k}$. Thus, plugging derived results into (8) and applying the property of the generalized Rayleigh quotient, we can easily obtain the closed-form results of $\mathbf{w}_{k}^{*}$ and $\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF},*}$ in Theorem 3.

### v2

@@ Appendix B Proof of Theorem 1
where $\sum_{\mathbf{m}}{}=\sum_{m_{1}=1}^{M}{\sum_{m_{2}=1}^{M}{\sum_{m_{3}=1}^{M}{\sum_{m_{4}=1}^{M}{}}}}$. Note that different phase-shifts with different $m$-subscript are i.i.d. and many cross-terms equal to 0 due to the existence of phase-shifts. By discussing all possible combinations of $m_{1}\sim m_{4}$, we can compute (43) into closed-form as shown in (10). As for $\mathbb{E}\{\|\mathbf{v}_{k}\|^{2}\}$, we have $\mathbb{E}\{\|\mathbf{v}_{k}\|^{2}\}=\mathbb{E}\{\widehat{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\mathbf{W}_{k}\widehat{\mathbf{g}}_{k}\}=\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\})=\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\overline{\mathbf{R}}_{k}).$ Finally, combining all results in this appendix, we finish the proof.
@@ Appendix C Proof of Theorem 2
We provide detailed proof steps of Theorem 2 in this appendix. We start from $\mathbb{E}\left\{\mathbf{b}_{kk}\right\}$ and we can easily compute the $m$-th element of it in closed-form as $\mathbb{E}\{\mathbf{v}_{mk}^{H}\mathbf{g}_{mk}\}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{R}}_{mk})$ based on the results in Appendix B. As for $[\mathbf{\Xi}_{kl}]_{nm}=\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}$, we can compute it in closed-form for all possible AP-UE combinations. For $m\neq n,l\notin\mathcal{P}_{k}$, we have $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\mathbb{E}\{\mathbf{g}_{ml}^{H}\mathbf{v}_{mk}\}\mathbb{E}\{\mathbf{v}_{nk}^{H}\mathbf{g}_{nl}\}=0$. For $m\neq n,l\in\mathcal{P}_{k}$, $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\mathrm{tr}(\mathbf{W}_{mk}\mathbf{B}_{mkl})\mathrm{tr}(\mathbf{W}_{nk}^{H}\mathbf{B}_{nlk})$, where $\mathbf{B}_{mkl}=\mathbb{E}\{\widehat{\mathbf{g}}_{mk}\mathbf{g}_{ml}^{H}\}=\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{ml}$ for $l\in\mathcal{P}_{k}\backslash\{k\}$ and $\overline{\mathbf{G}}_{mkk}+\sqrt{p_{k}p_{l}}\tau_{p}\check{\mathbf{R}}_{mk}\mathbf{\Psi}_{mk}^{-1}\check{\mathbf{R}}_{ml}$ for $l=k$. For $m=n,l\notin\mathcal{P}_{k}$, $\widehat{\mathbf{g}}_{mk}$ and $\mathbf{g}_{ml}$ are independent, so we can derive $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{G}}_{mll}\mathbf{W}_{mk}\overline{\mathbf{G}}_{mkk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\overline{\mathbf{G}}_{mkk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\overline{\mathbf{G}}_{mll}\widehat{\mathbf{R}}_{mk}\mathbf{W}_{mk})+\mathrm{tr}(\mathbf{W}_{mk}^{H}\check{\mathbf{R}}_{ml}\mathbf{W}_{mk}\widehat{\mathbf{R}}_{mk})$, by the method discussed in (38). As for $m=n,l\in\mathcal{P}_{k}$, $\widehat{\mathbf{g}}_{mk}$ and $\mathbf{g}_{ml}$ are not independent. Based on the similar method as (39), we can easily compute $\mathbb{E}\{(\mathbf{v}_{mk}^{H}\mathbf{g}_{ml})^{H}(\mathbf{v}_{nk}^{H}\mathbf{g}_{nl})\}=\varphi_{mkl}+\upsilon_{mkl}$. Moreover, we have $\mathbb{E}\{\|\mathbf{v}_{mk}\|^{2}\}=\mathrm{tr}(\mathbf{W}_{mk}^{H}\mathbf{W}_{mk}\overline{\mathbf{R}}_{mk})$. Based on all derived results, we can easily derive the closed-form results in Theorem 2.
@@ Appendix D Proof of Corollary 1
By applying (32), we can represent $\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF}}$ in (22) as (44) on the top of this page. By vectorizing $\mathbf{w}_{k}=\mathrm{vec}(\mathbf{W}_{k})\in\mathbb{C}^{M^{2}N^{2}}$, we can derive $\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})=\mathbf{w}_{k}^{H}\mathrm{vec}(\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\})$, applying (34) by letting $\mathbf{A}=\mathbf{W}_{k}$ and $\mathbf{B}=\mathbb{E}\{\mathbf{g}_{k}\widehat{\mathbf{g}}_{k}^{H}\}$, respectively. As for $\mathbb{E}\{\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{W}_{k}\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H})\}$, we can apply (36) to derive $\mathbb{E}\{\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{g}_{l}\mathbf{g}_{l}^{H}\mathbf{W}_{k}\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H})\}\!\!=\!\!\mathbf{w}_{k}^{H}\mathbb{E}\{[(\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H})^{T}\otimes(\mathbf{g}_{l}\mathbf{g}_{l}^{H})]\}\mathbf{w}_{k},$ by letting $\mathbf{A}=\mathbf{W}_{k}$, $\mathbf{B}=\mathbf{g}_{l}\mathbf{g}_{l}^{H}$, $\mathbf{C}=\mathbf{W}_{k}$, and $\mathbf{D}^{T}=\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}$, respectively. For $\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\})$, based on (35), we can obtain $\mathrm{tr}(\mathbf{W}_{k}^{H}\mathbf{W}_{k}\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\})=\mathbf{w}_{k}^{H}(\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\}^{T}\otimes\mathbf{I}_{MN})\mathbf{w}_{k}$, by letting $\mathbf{A}=\mathbf{W}_{k}$, $\mathbf{B}=\mathbf{I}_{MN}$, $\mathbf{C}=\mathbf{W}_{k}$, and $\mathbf{D}^{T}=\mathbb{E}\{\widehat{\mathbf{g}}_{k}\widehat{\mathbf{g}}_{k}^{H}\}$, respectively. Thus, by combining all these terms, we can construct (44) in (45) on the top of this page. Note that (45) is a standard Rayleigh quotient with respect to $\mathbf{w}_{k}$. Thus, by applying [5, Lemma B.10], we can derive the optimal $\mathbf{w}_{k}^{*}$, which can maximize (45), in (24), leading to the maximum SINR value in (23).
@@ Appendix E Proof of Theorem 3
We first reformulate terms of $\overline{\mathrm{SINR}}_{k}^{\mathrm{c},\mathrm{UatF}}$ given in (8). By applying (34), we have $\mathrm{tr}(\mathbf{W}_{k}^{H}\overline{\mathbf{R}}_{k})=\mathbf{w}_{k}^{H}\overline{\mathbf{r}}_{k}$. Moreover, by applying results in Appendix A, we can easily represent $\mu_{kl}$, $\varepsilon_{kl}$, and $\mathrm{tr}(\mathbf{W}_{k}\overline{\mathbf{R}}_{k}\mathbf{W}_{k}^{H})$ in (8) as $\mu_{kl}=\mathbf{w}_{k}^{H}(\overline{\mathbf{G}}_{kk}^{T}\otimes\overline{\mathbf{G}}_{ll})\mathbf{w}_{k}+\mathbf{w}_{k}^{H}(\overline{\mathbf{G}}_{kk}^{T}\otimes\check{\mathbf{R}}_{l})\mathbf{w}_{k}+\mathbf{w}_{k}^{H}(\widehat{\mathbf{R}}_{k}^{T}\otimes\overline{\mathbf{G}}_{ll})\mathbf{w}_{k}+\mathbf{w}_{k}^{H}(\widehat{\mathbf{R}}_{k}^{T}\otimes\check{\mathbf{R}}_{l})\mathbf{w}_{k}$, $\varepsilon_{kl}=p_{k}p_{l}\tau_{p}^{2}|\mathbf{w}_{k}^{H}\tilde{\mathbf{r}}_{lk}|^{2}$, and $\mathrm{tr}(\mathbf{W}_{k}\overline{\mathbf{R}}_{k}\mathbf{W}_{k}^{H})=\mathbf{w}_{k}^{H}(\overline{\mathbf{R}}_{k}^{T}\otimes\mathbf{I}_{MN})\mathbf{w}_{k}$, respectively. As for $\omega_{k}$ in (10), the first three terms can also be easily reformulated as $p_{k}\tau_{p}\mathbf{w}_{k}^{H}\overline{\mathbf{g}}_{kk}\tilde{\mathbf{r}}_{kk}^{H}\mathbf{w}_{k}+p_{k}\tau_{p}\mathbf{w}_{k}^{H}\tilde{\mathbf{r}}_{kk}\overline{\mathbf{g}}_{kk}^{H}\mathbf{w}_{k}-\mathbf{w}_{k}^{H}(\overline{\mathbf{G}}_{kk}^{T}\otimes\overline{\mathbf{G}}_{kk})\mathbf{w}_{k}$. However, as for the last three terms in (10), it is difficult to directly handle them since $\mathbf{W}_{k}$ has been divided into many block sub-matrices. To efficiently handle them, we focus on their original expectation form $\mathbb{E}\{\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\overline{\mathbf{g}}_{k}\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}\overline{\mathbf{g}}_{k}\}$. By applying (36), we have $\mathbb{E}\{\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}^{H}\overline{\mathbf{g}}_{k}\overline{\mathbf{g}}_{k}^{H}\mathbf{W}_{k}\overline{\mathbf{g}}_{k}\}=\mathbf{w}_{k}^{H}\underset{\mathbf{\Upsilon}_{k}}{\underbrace{\mathbb{E}\{(\overline{\mathbf{g}}_{k}\overline{\mathbf{g}}_{k}^{H})^{T}\otimes\overline{\mathbf{g}}_{k}\overline{\mathbf{g}}_{k}^{H}\}}}\mathbf{w}_{k}$. Then, we focus on the computation of $\mathbf{\Upsilon}_{k}\in\mathbb{C}^{M^{2}N^{2}\times M^{2}N^{2}}$, we define $\tilde{\mathbf{G}}_{k}=\overline{\mathbf{g}}_{k}\overline{\mathbf{g}}_{k}^{H}\in\mathbb{C}^{MN\times MN}$ and the $(d,c)$-th element of $\tilde{\mathbf{G}}_{k}$ is $\tilde{G}_{k,dc}=\overline{g}_{m_{1}k,n_{1}}\overline{g}_{m_{2}k,n_{2}}^{*}e^{j\theta_{m_{1}k}}e^{-j\theta_{m_{2}k}}$, where these variables are defined in Theorem 3. Then, by applying the characteristic of Kronecker-product, we can represent the $(c,d)$-th block sub-matrix of $\mathbf{\Upsilon}_{k}$ as $\mathbf{\Upsilon}_{k,cd}=\mathbb{E}\{\tilde{G}_{k,dc}\tilde{\mathbf{G}}_{k}\}\in\mathbb{C}^{MN\times MN}$. Finally, by plugging $\tilde{G}_{k,dc}$ into $\mathbf{\Upsilon}_{k,cd}$ and applying the features of random phase-shifts, we can easily compute $\mathbf{\Upsilon}_{k}$ in closed-form as $\overline{\mathbf{\Upsilon}}_{k}$. Thus, plugging derived results into (8) and applying the property of the generalized Rayleigh quotient, we can easily obtain the closed-form results of $\mathbf{w}_{k}^{*}$ and $\mathrm{SINR}_{k}^{\mathrm{c},\mathrm{UatF},*}$ in Theorem 3.